Exactly how many cubic inches of ice cream can an ice cream cone hold (within the cone) if it's height is 4 inches and the radius of it's base is 1.5 inches? Please use "pi" for the symbol in the problem. I am so completely lost, I have looked at book over and over again, I just don't understand...
Found 2 solutions by richwmiller, dylantc77:
Answer by richwmiller(17219)

What is the volume of a cone? h=4 and radius = 1.5

Answer by dylantc77(3)

v=pi*r^2*h
V = pi *4^2*1.5
V = Pi * 24
V = 75.3982